OpenStudy (anonymous):

cos(4x)+cos(2x)=2-2sin^2 (2x)-2sin^2(x)

OpenStudy (anonymous):

prove the identity

OpenStudy (freckles):

what have you tried

OpenStudy (anonymous):

i have been looking at all the identities but i dont know where to start

OpenStudy (anonymous):

ive tried sum difference identity but idk

OpenStudy (freckles):

I would take right and tried to show the left I think one of the key identities needed here is : \[2\cos^2(\theta)-1=\cos(2 \theta)\]

OpenStudy (freckles):

Or ... you could look at the identity \[1-2\sin^2(\theta)=\cos(2 \theta)\]

OpenStudy (freckles):

\[2-2\sin^2(2x)-2\sin^2(x) \\ =1+1-2\sin^2(2x)-2\sin^2(x) \\= 1-2\sin^2(2x)+1-2\sin^2(x)\]

OpenStudy (freckles):

If I go any further, I would have done the whole problem for you... Do you think you know where to go from here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks